OpenStudy (anonymous):

Anti derivative of sec^2(theta/2)

OpenStudy (anonymous):

\[\sec^2(\theta/2)\] How does this anti derivative = 2tan(theta/2), don't understand the "2" in the front part.

OpenStudy (psymon):

You're letting u = theta/2 u = theta/2 du = (1/2)dtheta dtheta = du/(1/2) = simply 2du \[\int\limits_{}^{}\sec^{2}(\theta /2)d \theta = \int\limits_{}^{}\sec^{2}u(2du)\]

OpenStudy (anonymous):

Ahh, nice, thanks Psymon :P, just stressing myself out, midterm tomorrow :X.

OpenStudy (psymon):

So ive heard, lol.